OpenStudy (anonymous):

does anyone know how to solve this long handed? log(base7)727 I don't have a graphing calculator and i need to be able to do it long handed for the test this week.

OpenStudy (amistre64):

7^x = 727 ; start out by seeing how close you can get with it: 7 = 7 7*7 = 49 7*7*7 = 343 7*7*7*7 = 2401; so its between 3 and 4

OpenStudy (amistre64):

change of base would help if you new a common base that 7 and 727 go into;

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thank you very much!

OpenStudy (amistre64):

yw, hope you do well :)